# Gravitational Lensing of Gravitational Waves from Astrophysical Sources: Theory, Detection, and Applications

### 07 May 2026 v1

@@ 3.1.2 Supermassive Binary Black Holes
The population synthesis model described above relies on the galaxy stellar mass function and merger rate with constraints from observations, which is effectively obtained only up to redshift $z\sim 5-6$ because of the paucity of high‑redshift observations. Extrapolating this to low-mass end of SMBBHs/MBBHs (e.g., $\lesssim 10^{5}-10^{6}M_{\odot}$) therefore may have substantial uncertainty. When considering merger rate of SMBBHs/MBBHs at low-mass end that may be detected by space-based GW detectors, more comprehensive approaches are required, which must follow galaxy assembly to much earlier times ($z\gtrsim 10-20$) and explicitly implement SMBH seeding; however seeding physics remains poorly understood and different seeding prescriptions produce large variations in predicted merger rates. For example, Barausse (2012) developed a semi-analytic model (SAM), based on the extended Press-Schechter (EPS) merger trees, to track mergers of host halos and galaxies across cosmic time and generate SMBBH catalogs for both light-seed ($\sim 10-100M_{\odot}$) and heavy-seed ($\sim 10^{4}-10^{6}M_{\odot}$) scenarios. Though, such SAM has since been refined to include more realistic physics, including accretion (Sesana et al. 2014), co-evolution with nuclear star cluster (NSC) (Antonini et al. 2015a, b), the time-delay between galaxy and SMBH mergers (Klein et al. 2016; Bonetti et al. 2018, 2019). kpc-scale wandering, and supernova feedback (Barausse et al. 2020), the SMBBH merger rate predictions remain highly model dependent and that progress requires both deeper high‑redshift observations and improved theoretical modeling of SMBH seeding and early growth. In addition, we also note other works in the literature (e.g., Jiang et al. 2019b; Wang et al. 2019; Salcido et al. 2016; Izquierdo-Villalba et al. 2019, 2020; Izquierdo-Villalba et al. 2022b) that directly adopted cosmological numerical simulation results to trace merger history of dark matter halos and galaxies, with implementations of SMBH seeding models and galaxy formation processes, to estimate SMBBH merger rates.
@@ 3.2 Lens Population
The propagation of GW can be deflected by any massive intervening objects/systems as discussed above in Section 2. In classical gravitational lensing studies of light propagation, the lensing effect is typically classified into three categories according to the nature of the lensing mass (Schneider et al. 2006): (1) strong lensing, caused by systems like galaxies and galaxy clusters; (2) microlensing, induced by objects such as individual stars, BHs, or planets; and (3) weak lensing, resulting from the cumulative influence of large scale structures in the universe. Similarly, GW lensing includes strong lensing by galaxies and galaxy clusters and microlensing by stars, black holes, or planets (e.g., Savastano et al. 2024; Zhang & Fan 2021; Shan et al. 2025b); for certain combinations of source frequency and lens Einstein radius, wave‑optics effects must be taken into account. However, weak lensing is sparingly studied for GWs and mentioned in the context of the stochastic background or as a noise term in distance measurement (e.g., Mpetha et al. 2024; Canevarolo & Chisari 2024). A newly recognized intermediate regime between wave- and geometric-optics — millilensing — features limited diffraction yet overlapping lensed GW signals whenever the inter‑image time delays are shorter than the waveform duration, yielding a frequency‑dependent interference (beating) pattern (e.g., Liu et al. 2023; Liu & Kim 2024; Liu et al. 2025). We refer readers to the above cited works for detailed treatments and limit our scope to lenses composed of galaxies, galaxy clusters, stars, and dark matter minihalos in both geometric and wave-optics regimes.
@@ 3.2.1 Galaxies
Strong lensing may be dominated by elliptical (early-type) galaxies (spiral galaxies may also contribute mildly‘, see Biesiada et al. (2014) ), for their larger mass $\sim 10^{8}-10^{12}M_{\odot}$ and compact density profile (with Sersic index $\sim 4$) (e.g., Oguri & Marshall 2010; Wong et al. 2017). The lens galaxies are therefore typically modelled by the Singular-Isothermal-Ellipsoidal (SIE) profile [^7] It is also frequently simply modelled as the Singular-Isothermal-Sphere (SIS) for rough estimations. (Kormann et al. 1994), of which the convergence in the Cartesian coordinates can be written as
@@
In the geometric-optics regime, the number of images due to strong lensing may depend on the lens density profile (e.g., Tessore & Metcalf 2015; Wierda et al. 2021). When assuming the SIE profile, the number of images with finite magnification factors are $2$ or $4$, and the fraction of $4$-image cases is much smaller comparing with that of $2$-image cases. However, adopting the more realistic lens models, such as Non-Singular Isothermal Ellipsoidal (NSIE) or Double Elliptic Power-Law (DEPL) - motivated by the existence of core-like structure in early-type galaxies (Byun et al. 1996; Lauer et al. 1997)- can produce additional images (including a central image) that do not appear in contrast to the simpler SIE model. The central image, arising from the core-like structure of the lens galaxy, has traditionally received limited attention for several reasons. First, it is normally much fainter than the other images due to its low (de)magnification factor, typically around $\mu\sim 0.01$. Second, its location is close to the optical axis of the lens system, making it susceptible to strong contamination from the lens galaxy’s own emission. Nonetheless, incorporating a core-like structure into the lens galaxy is a common practice when modeling member galaxies in cluster lensing systems (e.g., Xie et al. 2024). Notably, two radio-detected lensed Quasars, with three (Winn et al. 2004) and five images (Inada et al. 2005), respectively, show central images that are plausibly attributable to core-like mass distribution in their lensing galaxies. When considering lensed GW events, the above two difficulties in detecting the central image, if any, may be naturally overcome. First, the S/N of the central image may be only about ten times fainter than other images as its signal $\propto\sqrt{\mu}$, which may be still detectable by the third generation GW detector despite that it is relatively weak comparing with other images. Second, the GW signal can get through the lens galaxy without any contamination though it may be affected slightly by the micro-lensing effect. It has been shown that the detection of the central image is probable (Chen et al. 2025a).
@@ 3.2.2 Galaxy clusters
Modeling galaxy clusters is challenging because of their complex structures-multiple clumpy dark matter halos and numerous dynamically perturbed substructures. However, Chen et al. (2024d) found that the overall detection rate of the cluster-lensed GW events, whether estimated from mock clusters in the Synthetic Sky Catalog for Dark Energy Science with LSST (cosmoDC2) (e.g., Korytov et al. 2019; LSST Dark Energy Science Collaboration 2012) or derived using an eccentric Navarro–Frenk–White (eNFW; Navarro et al. 1997; Wright & Brainerd 2000; Golse & Kneib 2002) dark matter halo plus a brightest central galaxy (BCG) with SIS profile, are in agreement with each other. This consistency arises because small-scale substructures introduce only minor perturbation to the combined lensing potential of the BCG and the dark matter halo. Therefore, adopting the eNFW+SIS profile offers a plausible and efficient approximation for estimating the detectability of cluster-lensed GW events. On the one hand, the density profile of the NFW profile is
@@
where $\sigma_{\rm v}$ is the velocity dispersion of the BCG. The BCG can be linked to the main dark matter halo by adopting the extended empirical $M_{\ast,\rm BCG}-M_{\ell}$ relationship (for example, see Erfanianfar et al. 2019). Then the velocity dispersion $\sigma_{\rm v}$ of the BCG can be estimated via the BCG mass according to observations (for example, see Hyde & Bernardi 2009). Notably, unlike the case for galaxy lensing, the above joint lens model may only be solved numerically rather than analytically, for their complex math structures.
In the above eNFW+SIS model for cluster lensing, the key parameter is the dark matter halo mass $M_{\ell}$. The distribution of the lens cluster population is given by the cosmic halo mass function (HMF). For example, a widely adopted general expression is given by Sheth & Tormen (1999)
@@
where $\rho_{0}$ is the mean mass density of the universe, $\sigma(M)$ is the variance of density fluctuations in the universe on a scale corresponding to mass $M$, and $\nu$ is the dimensionless parameter defined as $\nu=\delta_{\rm c}/\sigma(M)$, in which $\delta_{\rm c}$ is the critical density threshold for collapse (approximately $\delta_{\rm c}\approx 1.686(1+0.123\log_{10}{}\Omega_{\Lambda})$ depending on the dark energy density $\Omega_{\Lambda}$ (Kitayama & Suto 1996)). The fitting parameters $a$ and $p$ are suggested to be $(a,p)=(0.75,0.3)$ by Sheth & Tormen (1999) and notably when $(a,p)=(1,0)$, it restores the original Press-Schechter function (Press & Schechter 1974). We refer the interested readers to Python package hmf to generate HMF in various forms (Murray et al. 2013). The HMF behaves approximately as a power-law over $10^{13}-10^{14}M_{\odot}$ and declines rapidly above $10^{14}M_{\odot}$, implying that strong lensing by galaxy clusters is dominated by the low mass end of the cluster population.
@@ 3.2.3 Stars
Stars within galaxies can also play a role in gravitational lensing as the intervening object, which corresponds to the microlensing regime (e.g., Paczynski 1986; Mao & Paczynski 1991; Alcock et al. 1997; Beaulieu et al. 2006; Mao 2012; Gaudi 2012; Yan et al. 2014; Mishra et al. 2021; Shan et al. 2021, 2025b). As for the GW lensing, the condition $\lambda_{\rm GW}/\lambda_{E}\gtrsim 1$ is always satisfied in this case. Therefore, the gravitational lensing of GW by stars should be considered within the wave-optics regime. Due to the small size of stars, we often model them as superposition of $N$ point sources, of which the lensing potential is
@@ 3.2.4 Primordial Black holes
The primordial black holes (PBH) are originated from the gravitational collapse of the overdense regions in the early universe (Zel’dovich & Novikov 1967; Hawking 1971), which is a fancy candidate for dark matter. The optical microlensing effect of PBHs has been widely used to constrain the mass of PBHs. In principle, PBHs can also serve as lenses for the GW signals (e.g., Cai et al. 2023; Urrutia et al. 2023). Here we divide the PBH lensing into two categories. The first one is for bare PBHs, which can be modeled as simple point mass lenses. The second one is for dressed PBHs, considering the minihalo around each PBH caused by accretion of surrounding dark matter (e.g., Bertschinger 1985; Eroshenko 2016). In the latter case, the density profiles of the surrounding minihalos are often parameterized as  [^8] The power law index 9/4-9/4 is originally obtained by Bertschinger (1985) assuming spherical collapse in Einstein-de Sitter Universe without central BHs, which may not be the case for dressed PBHs. Analogous to the collisionless particles around the BH, the density profile of dark matter halo may be sketched by the Bahcall-Wolf cusp with index of 7/4-7/4 derived from galactic dynamics (Bahcall & Wolf 1976) rather than 9/4-9/4. (e.g., Adamek et al. 2019; Cai et al. 2023):
@@
where $M_{\rm halo}$ is the total mass of the halo and $R_{\rm halo}$ is the cutoff radius of the halo density profile. These two quantities are directly related to the PBH mass and the fraction of dark matter in the forms of PBH, i.e., $m_{\rm PBH}$ and $f_{\rm PBH}$. A simple argument is that $M_{\rm halo}$ is limited by the available amount of particle dark matter in the universe, which implies that $f_{\rm PBH}M_{\rm halo}\leq(1-f_{\rm PBH})m_{\rm PBH}$. The equality is only satisfied in the idealized case when all of the dark matter is contained in PBH minihalos. One may also expect a stronger condition that the minihalo growth is stopped once the PBHs are absorbed into nonlinear dark matter structures, which imposes a limitation such as $1+z_{\rm NL}\sim 5500f_{\rm PBH}$ (e.g., Inman & Ali-Haïmoud 2019), where $z_{\rm NL}$ is the redshift for the formation of nonlinear dark matter structures. Then, the size of the dressed PBH minihalo $R_{\rm halo}$ can be defined by the radius at which the density of the minihalos matches the ambient dark matter density, i.e., $\rho_{\rm R_{\rm halo}}\sim\rho_{\rm DM}(z_{\rm NL})$. Therefore, we have
@@ 3.2.5 Mini Dark Matter Halo
The GW emitted by CBC events can also be gravitationally lensed by mini dark matter halos with mass $\sim 10^{3}-10^{6}M_{\odot}$ (hereafter minihalos), with significant cross-section due to wave effect (also diffraction lensing), compared with EM signals. These systems can be viewed as ideal probe to the nature of dark matter, as they are not polluted by sophisticated baryon processes (e.g., Guo & Lu 2022; Tambalo et al. 2023; Jana et al. 2024; Cheung et al. 2024; Singh et al. 2025). For different dark matter models, the HMF for minihalos given by Equation (59) may vary. For example, the warm dark matter (WDM) model tends to predict less minihalos at the lower mass end compared with cold dark matter (CDM) model, depending on the particle masses. More importantly, the density profiles of minihalos produced by different dark matter models may be significantly different from each other, especially at the halo centers. Here we make a comparison between the CDM and self-interaction dark matter (SIDM) model as an example, referring the readers to Vegetti et al. (2024) for extensive discussion on other dark matter models, including WDM and fuzzy dark matter.

### 16 May 2026 v2

@@ 3.1.2 Supermassive Binary Black Holes
The population synthesis model described above relies on the galaxy stellar mass function and merger rate with constraints from observations, which is effectively obtained only up to redshift $z\sim 5-6$ because of the paucity of high‑redshift observations. Extrapolating this to low-mass end of SMBBHs/MBBHs (e.g., $\lesssim 10^{5}-10^{6}M_{\odot}$) therefore may have substantial uncertainty. When considering merger rate of SMBBHs/MBBHs at low-mass end that may be detected by space-based GW detectors, more comprehensive approaches are required, which must follow galaxy assembly to much earlier times ($z\gtrsim 10-20$) and explicitly implement SMBH seeding; however seeding physics remains poorly understood and different seeding prescriptions produce large variations in predicted merger rates. For example, Barausse (2012) developed a semi-analytic model (SAM), based on the extended Press-Schechter (EPS) merger trees, to track mergers of host halos and galaxies across cosmic time and generate SMBBH catalogs for both light-seed ($\sim 10-100M_{\odot}$) and heavy-seed ($\sim 10^{4}-10^{6}M_{\odot}$) scenarios. Though, such SAM has since been refined to include more realistic physics, including accretion (Sesana et al. 2014), co-evolution with nuclear star cluster (NSC) (Antonini et al. 2015b, a), the time-delay between galaxy and SMBH mergers (Klein et al. 2016; Bonetti et al. 2018, 2019). kpc-scale wandering, and supernova feedback (Barausse et al. 2020), the SMBBH merger rate predictions remain highly model dependent and that progress requires both deeper high‑redshift observations and improved theoretical modeling of SMBH seeding and early growth. In addition, we also note other works in the literature (e.g., Jiang et al. 2019b; Wang et al. 2019; Salcido et al. 2016; Izquierdo-Villalba et al. 2019, 2020, 2022b) that directly adopted cosmological numerical simulation results to trace merger history of dark matter halos and galaxies, with implementations of SMBH seeding models and galaxy formation processes, to estimate SMBBH merger rates.
@@ 3.2 Lens Population
The propagation of GW can be deflected by any massive intervening objects/systems as discussed above in Section 2. In classical gravitational lensing studies of light propagation, the lensing effect is typically classified into three categories according to the nature of the lensing mass (Schneider et al. 2006): (1) strong lensing, caused by systems like galaxies and galaxy clusters; (2) microlensing, induced by objects such as individual stars, BHs, or planets; and (3) weak lensing, resulting from the cumulative influence of large scale structures in the universe. Similarly, GW lensing includes strong lensing by galaxies and galaxy clusters and microlensing by stars, black holes, or planets (e.g., Savastano et al. 2024; Zhang and Fan 2021; Shan et al. 2025b); for certain combinations of source frequency and lens Einstein radius, wave‑optics effects must be taken into account. However, weak lensing is sparingly studied for GWs and mentioned in the context of the stochastic background or as a noise term in distance measurement (e.g., Mpetha et al. 2024; Canevarolo and Chisari 2024). A newly recognized intermediate regime between wave- and geometric-optics — millilensing — features limited diffraction yet overlapping lensed GW signals whenever the inter‑image time delays are shorter than the waveform duration, yielding a frequency‑dependent interference (beating) pattern (e.g., Liu et al. 2023; Liu and Kim 2024; Liu et al. 2025). We refer readers to the above cited works for detailed treatments and limit our scope to lenses composed of galaxies, galaxy clusters, stars, and dark matter minihalos in both geometric and wave-optics regimes.
@@ 3.2.1 Galaxies
Strong lensing may be dominated by elliptical (early-type) galaxies (spiral galaxies may also contribute mildly‘, see Biesiada et al. (2014) ), for their larger mass $\sim 10^{8}-10^{12}M_{\odot}$ and compact density profile (with Sersic index $\sim 4$) (e.g., Oguri and Marshall 2010; Wong et al. 2017). The lens galaxies are therefore typically modelled by the Singular-Isothermal-Ellipsoidal (SIE) profile [^9] It is also frequently simply modelled as the Singular-Isothermal-Sphere (SIS) for rough estimations. (Kormann et al. 1994), of which the convergence in the Cartesian coordinates can be written as
@@
In the geometric-optics regime, the number of images due to strong lensing may depend on the lens density profile (e.g., Tessore and Metcalf 2015; Wierda et al. 2021). When assuming the SIE profile, the number of images with finite magnification factors are $2$ or $4$, and the fraction of $4$-image cases is much smaller comparing with that of $2$-image cases. However, adopting the more realistic lens models, such as Non-Singular Isothermal Ellipsoidal (NSIE) or Double Elliptic Power-Law (DEPL) - motivated by the existence of core-like structure in early-type galaxies (Byun et al. 1996; Lauer et al. 1997)- can produce additional images (including a central image) that do not appear in contrast to the simpler SIE model. The central image, arising from the core-like structure of the lens galaxy, has traditionally received limited attention for several reasons. First, it is normally much fainter than the other images due to its low (de)magnification factor, typically around $\mu\sim 0.01$. Second, its location is close to the optical axis of the lens system, making it susceptible to strong contamination from the lens galaxy’s own emission. Nonetheless, incorporating a core-like structure into the lens galaxy is a common practice when modeling member galaxies in cluster lensing systems (e.g., Xie et al. 2024). Notably, two radio-detected lensed Quasars, with three (Winn et al. 2004) and five images (Inada et al. 2005), respectively, show central images that are plausibly attributable to core-like mass distribution in their lensing galaxies. When considering lensed GW events, the above two difficulties in detecting the central image, if any, may be naturally overcome. First, the S/N of the central image may be only about ten times fainter than other images as its signal $\propto\sqrt{\mu}$, which may be still detectable by the third generation GW detector despite that it is relatively weak comparing with other images. Second, the GW signal can get through the lens galaxy without any contamination though it may be affected slightly by the micro-lensing effect. It has been shown that the detection of the central image is probable (Chen et al. 2025b).
@@ 3.2.2 Galaxy clusters
Modeling galaxy clusters is challenging because of their complex structures-multiple clumpy dark matter halos and numerous dynamically perturbed substructures. However, Chen et al. (2024d) found that the overall detection rate of the cluster-lensed GW events, whether estimated from mock clusters in the Synthetic Sky Catalog for Dark Energy Science with LSST (cosmoDC2) (e.g., Korytov et al. 2019; LSST Dark Energy Science Collaboration 2012) or derived using an eccentric Navarro–Frenk–White (eNFW; Navarro et al. 1997; Wright and Brainerd 2000; Golse and Kneib 2002) dark matter halo plus a brightest central galaxy (BCG) with SIS profile, are in agreement with each other. This consistency arises because small-scale substructures introduce only minor perturbation to the combined lensing potential of the BCG and the dark matter halo. Therefore, adopting the eNFW+SIS profile offers a plausible and efficient approximation for estimating the detectability of cluster-lensed GW events. On the one hand, the density profile of the NFW profile is
@@
where $\sigma_{\rm v}$ is the velocity dispersion of the BCG. The BCG can be linked to the main dark matter halo by adopting the extended empirical $M_{\ast,\rm BCG}-M_{\ell}$ relationship (for example, see Erfanianfar et al. 2019). Then the velocity dispersion $\sigma_{\rm v}$ of the BCG can be estimated via the BCG mass according to observations (for example, see Hyde and Bernardi 2009). Notably, unlike the case for galaxy lensing, the above joint lens model may only be solved numerically rather than analytically, for their complex math structures.
In the above eNFW+SIS model for cluster lensing, the key parameter is the dark matter halo mass $M_{\ell}$. The distribution of the lens cluster population is given by the cosmic halo mass function (HMF). For example, a widely adopted general expression is given by Sheth and Tormen (1999)
@@
where $\rho_{0}$ is the mean mass density of the universe, $\sigma(M)$ is the variance of density fluctuations in the universe on a scale corresponding to mass $M$, and $\nu$ is the dimensionless parameter defined as $\nu=\delta_{\rm c}/\sigma(M)$, in which $\delta_{\rm c}$ is the critical density threshold for collapse (approximately $\delta_{\rm c}\approx 1.686(1+0.123\log_{10}{}\Omega_{\Lambda})$ depending on the dark energy density $\Omega_{\Lambda}$ (Kitayama and Suto 1996)). The fitting parameters $a$ and $p$ are suggested to be $(a,p)=(0.75,0.3)$ by Sheth and Tormen (1999) and notably when $(a,p)=(1,0)$, it restores the original Press-Schechter function (Press and Schechter 1974). We refer the interested readers to Python package hmf to generate HMF in various forms (Murray et al. 2013). The HMF behaves approximately as a power-law over $10^{13}-10^{14}M_{\odot}$ and declines rapidly above $10^{14}M_{\odot}$, implying that strong lensing by galaxy clusters is dominated by the low mass end of the cluster population.
@@ 3.2.3 Stars
Stars within galaxies can also play a role in gravitational lensing as the intervening object, which corresponds to the microlensing regime (e.g., Paczynski 1986; Mao and Paczynski 1991; Alcock et al. 1997; Beaulieu et al. 2006; Mao 2012; Gaudi 2012; Yan et al. 2014; Mishra et al. 2021; Shan et al. 2021, 2025b). As for the GW lensing, the condition $\lambda_{\rm GW}/\lambda_{E}\gtrsim 1$ is always satisfied in this case. Therefore, the gravitational lensing of GW by stars should be considered within the wave-optics regime. Due to the small size of stars, we often model them as superposition of $N$ point sources, of which the lensing potential is
@@ 3.2.4 Primordial Black holes
The primordial black holes (PBH) are originated from the gravitational collapse of the overdense regions in the early universe (Zel’dovich and Novikov 1967; Hawking 1971), which is a fancy candidate for dark matter. The optical microlensing effect of PBHs has been widely used to constrain the mass of PBHs. In principle, PBHs can also serve as lenses for the GW signals (e.g., Cai et al. 2023; Urrutia et al. 2023). Here we divide the PBH lensing into two categories. The first one is for bare PBHs, which can be modeled as simple point mass lenses (e.g., Nakamura 1998; Deka et al. 2025). The second one is for dressed PBHs, considering the minihalo around each PBH caused by accretion of surrounding dark matter (e.g., Bertschinger 1985; Eroshenko 2016). In the latter case, the density profiles of the surrounding minihalos are often parameterized as  [^10] The power law index 9/4-9/4 is originally obtained by Bertschinger (1985) assuming spherical collapse in Einstein-de Sitter Universe without central BHs, which may not be the case for dressed PBHs. Analogous to the collisionless particles around the BH, the density profile of dark matter halo may be sketched by the Bahcall-Wolf cusp with index of 7/4-7/4 derived from galactic dynamics (Bahcall and Wolf 1976) rather than 9/4-9/4. (e.g., Adamek et al. 2019; Cai et al. 2023):
@@
where $M_{\rm halo}$ is the total mass of the halo and $R_{\rm halo}$ is the cutoff radius of the halo density profile. These two quantities are directly related to the PBH mass and the fraction of dark matter in the forms of PBH, i.e., $m_{\rm PBH}$ and $f_{\rm PBH}$. A simple argument is that $M_{\rm halo}$ is limited by the available amount of particle dark matter in the universe, which implies that $f_{\rm PBH}M_{\rm halo}\leq(1-f_{\rm PBH})m_{\rm PBH}$. The equality is only satisfied in the idealized case when all of the dark matter is contained in PBH minihalos. One may also expect a stronger condition that the minihalo growth is stopped once the PBHs are absorbed into nonlinear dark matter structures, which imposes a limitation such as $1+z_{\rm NL}\sim 5500f_{\rm PBH}$ (e.g., Inman and Ali-Haïmoud 2019), where $z_{\rm NL}$ is the redshift for the formation of nonlinear dark matter structures. Then, the size of the dressed PBH minihalo $R_{\rm halo}$ can be defined by the radius at which the density of the minihalos matches the ambient dark matter density, i.e., $\rho_{\rm R_{\rm halo}}\sim\rho_{\rm DM}(z_{\rm NL})$. Therefore, we have
@@ 3.2.5 Mini Dark Matter Halo
The GW emitted by CBC events can also be gravitationally lensed by mini dark matter halos with mass $\sim 10^{3}-10^{6}M_{\odot}$ (hereafter minihalos), with significant cross-section due to wave effect (also diffraction lensing), compared with EM signals. These systems can be viewed as ideal probe to the nature of dark matter, as they are not polluted by sophisticated baryon processes (e.g., Guo and Lu 2022; Tambalo et al. 2023; Jana et al. 2024; Cheung et al. 2024; Singh et al. 2025). For different dark matter models, the HMF for minihalos given by Equation (59) may vary. For example, the warm dark matter (WDM) model tends to predict less minihalos at the lower mass end compared with cold dark matter (CDM) model, depending on the particle masses. More importantly, the density profiles of minihalos produced by different dark matter models may be significantly different from each other, especially at the halo centers. Here we make a comparison between the CDM and self-interaction dark matter (SIDM) model as an example, referring the readers to Vegetti et al. (2024) for extensive discussion on other dark matter models, including WDM and fuzzy dark matter.